-3/2x + 11/4=-31/4

Answer by profemmanuel2q@yahoo.com(15)

-3/2x + 11/4=-31/4
-3/2x = -31/4 - 11/4 ( Solve -31/4 - 11/4 by finding the LCM)So you will get
-3/2x = -42/4 ( You just cross multiply)
2x x -42 = -3 x 4
-84x = -12 ( Divide both sides by - 84
-84x/-84 = -12/-84 ( Just conceal do the cancealation
So you have x = 1/7